1.Convert the polar coordinates (-8, ) into rectangular coordinates.

2.Convert the rectangular coordinates (3, -3) into polar coordinates with r > 0 and 0 ≤ θ < 2π.

3.Convert the rectangular equation x2 + y2 = 100 into a polar equation that expresses r in terms of θ.

4.Convert the polar equation 4r cos θ + r sin θ = 8 into a rectangular equation that expresses y in terms of x.




1
Expert's answer
2021-11-29T08:35:00-0500

"x=rcos\theta,y=rsin\theta"


1.:

"x=8cos\theta,y=8sin\theta"


2.

"3=rcos\theta,-3=rsin\theta"

"tan\theta=-1"

since "x>0,y<0"

"\theta=7\pi/4"

"r=3/cos(7\pi/4)=3\sqrt 2"


3.

"(rcos\theta)^2+(rsin\theta)^2=100"

"r^2=100"

"r=10"


4.

"4x+y=8"

"y=8-4x"